OpenStudy (anonymous):

Four more than twice a number is seven more than the number. What is the number?

OpenStudy (anonymous):

Let the number be X. 4 + 2X = X + 7

OpenStudy (anonymous):

Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1x' to each side of the equation. 4 + 2x + -1x = 7 + x + -1x Combine like terms: 2x + -1x = 1x 4 + 1x = 7 + x + -1x Combine like terms: x + -1x = 0 4 + 1x = 7 + 0 4 + 1x = 7 Add '-4' to each side of the equation. 4 + -4 + 1x = 7 + -4 Combine like terms: 4 + -4 = 0 0 + 1x = 7 + -4 1x = 7 + -4 Combine like terms: 7 + -4 = 3 1x = 3 Divide each side by '1'. x = 3 Simplifying x = 3